OpenStudy (anonymous):

Please Help Solve the equation and check the solution -7(t-5)-(t+4)=11 The solution is t=

OpenStudy (anonymous):

distribute -7t+35-t+4=11 add like terms -8t+39=11 -39 -39 -8t=-28 now divdide t=-28/-8 or 3.5

OpenStudy (anonymous):

so answer is 3.5 ?

jhonyy9 (jhonyy9):

@callie2240 there not will be +4